# Recent quantum runtime (dis)advantages

### v2

@@ Acknowledgments
We begin with providing a more detailed description of the HUBO instances used in Section IV of the main text. These instances are designed to be challenging, yet well suited for the IBM quantum computers, thus the starting point in their construction is the heavy-hex graph $C_{0}$ of IBM Heron architecture [65]. Two conflict graphs are then constructed: $C_{0}^{(2)}$ and $C_{0}^{(3)}$, with vertices corresponding to the edges of $C_{0}$ in the case of $C_{0}^{(2)}$, and to the triangles and three-qubit paths in the case of $C_{0}^{(3)}$. Edges in the conflict graphs connect vertices that share a vertex in $C_{0}$, which means that the operations on qubits encoding these interactions cannot be executed simultaneously. By applying graph coloring to the conflict graphs, one obtains partitions of the edges of $C_{0}$ into sets of non-conflicting edges, $P_{2q}=\{P^{(1)}_{2q},P^{(2)}_{2q},\ldots,P^{(M_{2})}_{2q}\}$ and $P_{3q}=\{P^{(1)}_{3q},P^{(2)}_{3q},\ldots,P^{(M_{3})}_{3q}\}$, where $M_{2}$ and $M_{3}$ are the number of colors used in the coloring of $C_{0}^{(2)}$ and $C_{0}^{(3)}$, respectively. Each set $P^{(i)}_{2q}$ and $P^{(i)}_{3q}$ contain two- and three-body interactions that can be executed in parallel on the quantum hardware.
@@
which introduces one auxiliary spin $s_{\rm aux}$ per third-order term, and preserves the spectrum of the original problem. In the GitHub repository [60], we provide both HUBO instances and their QUBO counterparts after the reduction. We also note, that SA can be extended to directly handle higher-order interactions fairly easily, and while in principle SBM can be too [67], it is much more involved. Thus, we only consider the HUBO-native version of SA, which we discuss in greater detail in App. A4.
@@
FIG. A1. : Success fraction as a function of instance size, for $50$ considered instances of each type. Panels (a) and (d) show results for SBM, (b) and (e) for SA in QUBO form, and (c) and (f) for SA in HUBO form. The top row (a-c) corresponds to instances with $S_{2q}=1$, $S_{3q}=4$, Cauchy distributed couplings, while the bottom row (d-f) to instances with $S_{2q}=1$, $S_{3q}=6$, symmetrized Pareto distributed couplings. These results show that, when it comes pure solution quality, these instances are challenging for Simulated Bifurcation. However, the success fraction can be significantly improved by increasing the number of replicas evolved in parallel, for which SBM is particularly well suited. SA, both in QUBO and HUBO form, performs significantly better, with the QUBO-native version being the best for Cauchy instances, and HUBO-native for Pareto instances.
@@
We start by analyzing the success fraction, defined as the fraction of instances for which a given algorithm found solution better than or equal to a certain threshold. The results are shown in Fig. A3.2, for three selected thresholds, $\mathcal{R}\in\{0.90,0.95,0.99\}$, and as a function of instance size $N$. We see a first hint that these instances are indeed challenging, since the success fraction decreases rapidly with both instance size and the threshold approximation ratio. Interestingly, Simulated Bifurcation seems to struggle much more than Simulated Annealing, at least when it comes to pure solution quality. However, when discussing approximate optimization algorithms, one should always be mindful of the time-quality trade-off. Let us now consider the runtime of the algorithms, to see a different picture emerge. In Fig. A3.2, we show the approximation ratio $\mathcal{R}$, averaged over $50$ couplings realizations and $10$ (55) independent runs of SBM (SA), as a function of runtime, for each considered instance type and size. Immediately, we can notice that while QUBO SA reaches better solutions eventually, SBM is able to reach good solutions close to an order of magnitude faster. Similar conclusions can be drawn when comparing SBM to HUBO SA in the case of Cauchy instances, while the Pareto instances seem to be “easy” for HUBO SA. Nevertheless, even in this case, the performance of SBM is still respectable, especially when considering the potential for further improvent via extending it to natively handle higher-order interactions [67].
@@
Crucially, this is a result of a statistically sound benchmarking procedure, being aware of, and taking into account the variability of instances due to random coupling selection. The dangers of cherry-picking results, which were discussed in the main text, here are clearly visible in the insets of Fig. A3.2, which show the rather broad distribution of best approximation ratios obtained for each instance type and size. To stress it again here, in Fig. A3.2 we reproduce Fig. IV from the main text, but with the addition of results from an optimized, GPU-based implementation of HUBO SA. This shows that not only the choice of baseline algorithm matters, but also the specific implementation can influence the results, and one should always strive to use the best available version of a given algorithm.
@@
FIG. A2. : Approximation ratio $\mathcal{R}$ as a function of runtime, averaged over $50$ instances and $10$ (55) independent runs of SBM (SA), for each instance type and size. Shaded area corresponds to one standard deviation, and insets show the distribution of $\mathcal{R}$ obtained, by aggregating best values for a given instance type and size. We can clearly see the biggest advantage of SBM over SA – its ability to efficiently evolve many replicas simultaneously, while keeping the runtime low. In terms of quality-runtime trade-off, it holds its ground even againts the HUBO SA variant, which does not suffer overheads from rank reduction. In particular, SBM is able to, on average, reach solutions of reasonable quality, $\mathcal{R}\approx 0.95$ for Cauchy instances and $\mathcal{R}\approx 0.90$ for Pareto instances, close to an order of magnitude faster than SA, except for large Pareto instances, which seems to be “easy” for the HUBO SA. This is promising, given that it is possible to extend SBM so that it can handle higher-order interactions natively [67]. Finally, insets reveal broad distribution of approximation ratios over random coupling selection for a fixed topology, which further highlights the importance of conducting a proper statistical analysis before drawing any conclusions.
@@
Appendix A4 Simulated Annealing for QUBO add HUBO
@@
  $num\_passses\in\mathbb{N}$ – number of passes over variables per temperature,
@@
  for $pass\in\{1,\,\ldots,\,num\_passses\}$ do
@@
All classical benchmarks in this work were performed on a workstation equipped with dual Intel(R) Xeon(R) Platinum 8462Y+ CPUs, $4$ NVIDIA H100 SXM GPUs, and 11TB of RAM. All results were obtained using a single GPU, unless stated otherwisee.

### v3

@@ Acknowledgments
We begin with providing a more detailed description of the HUBO instances used In Sec. IV of the main text. These instances are designed to be challenging, yet well suited for the IBM quantum computers, thus the starting point in their construction is the heavy-hex graph $C_{0}$ of IBM Heron architecture [69]. Two conflict graphs are then constructed: $C_{0}^{(2)}$ and $C_{0}^{(3)}$, with vertices corresponding to the edges of $C_{0}$ in the case of $C_{0}^{(2)}$, and to the triangles and three-qubit paths in the case of $C_{0}^{(3)}$. Edges in the conflict graphs connect vertices that share a vertex in $C_{0}$, which means that the operations on qubits encoding these interactions cannot be executed simultaneously. By applying graph coloring to the conflict graphs, one obtains partitions of the edges of $C_{0}$ into sets of non-conflicting edges, $P_{2q}=\{P^{(1)}_{2q},P^{(2)}_{2q},\ldots,P^{(M_{2})}_{2q}\}$ and $P_{3q}=\{P^{(1)}_{3q},P^{(2)}_{3q},\ldots,P^{(M_{3})}_{3q}\}$, where $M_{2}$ and $M_{3}$ are the number of colors used in the coloring of $C_{0}^{(2)}$ and $C_{0}^{(3)}$, respectively. Each set $P^{(i)}_{2q}$ and $P^{(i)}_{3q}$ contain two- and three-body interactions that can be executed in parallel on the quantum hardware.
@@
which introduces one auxiliary spin $s_{\rm aux}$ per third-order term, and preserves the spectrum of the original problem. In the GitHub repository [62], we provide both HUBO instances and their QUBO counterparts after the reduction. We also note that SA can be extended to directly handle higher-order interactions fairly easily, and while in principle SBM can be too [71], it is much more involved. Thus, we only consider the HUBO-native version of SA, which we discuss in greater detail in App. A4.
@@
FIG. A1. : Success fraction as a function of instance size, for $50$ considered instances of each type. Panels (a) and (d) show results for SBM, (b) and (e) for SA in QUBO form, and (c) and (f) for SA in HUBO form. The top row (a-c) corresponds to instances with $S_{2q}=1$, $S_{3q}=4$, Cauchy distributed couplings, while the bottom row (d-f) to instances with $S_{2q}=1$, $S_{3q}=6$, symmetrized Pareto distributed couplings. These results show that, when it comes to pure solution quality, these instances are challenging for Simulated Bifurcation. However, the success fraction can be significantly improved by increasing the number of replicas evolved in parallel, for which SBM is particularly well suited. SA, both in QUBO and HUBO form, performs significantly better, with the QUBO-native version being the best for Cauchy instances, and HUBO-native for Pareto instances.
@@
We start by analyzing the success fraction, defined as the fraction of instances for which a given algorithm found solution better than or equal to a certain threshold. The results are shown in Fig. A3.2, for three selected thresholds, $\mathcal{R}\in\{0.90,0.95,0.99\}$, and as a function of instance size $N$. We see a first hint that these instances are indeed challenging, since the success fraction decreases rapidly with both instance size and the threshold approximation ratio. Interestingly, Simulated Bifurcation seems to struggle much more than Simulated Annealing, at least when it comes to pure solution quality. However, when discussing approximate optimization algorithms, one should always be mindful of the time-quality trade-off. Let us now consider the runtime of the algorithms, to see a different picture emerge. In Fig. A3.2, we show the approximation ratio $\mathcal{R}$, averaged over $50$ coupling realizations and $10$ (55) independent runs of SBM (SA), as a function of runtime, for each considered instance type and size. Immediately, we can notice that while QUBO SA reaches better solutions eventually, SBM is able to reach good solutions close to an order of magnitude faster. Similar conclusions can be drawn when comparing SBM to HUBO SA in the case of Cauchy instances, while the Pareto instances seem to be “easy” for HUBO SA. Nevertheless, even in this case, the performance of SBM is still respectable, especially when considering the potential for further improvement via extending it to natively handle higher-order interactions [71].
@@
Crucially, this is a result of a statistically sound benchmarking procedure, being aware of, and taking into account the variability of instances due to random coupling selection. The impact of selective reporting, which were discussed in the main text, here are clearly visible in the insets of Fig. A3.2, which show the rather broad distribution of best approximation ratios obtained for each instance type and size. To stress it again here, in Fig. A3.2 we reproduce Fig. IV from the main text, but with the addition of results from an optimized, GPU-based implementation of HUBO SA. This shows that not only the choice of baseline algorithm matters, but also the specific implementation can influence the results, and one should always strive to use the best available version of a given algorithm.
@@
FIG. A2. : Approximation ratio $\mathcal{R}$ as a function of runtime, averaged over $50$ instances and $10$ (55) independent runs of SBM (SA), for each instance type and size. Shaded area corresponds to one standard deviation, and insets show the distribution of $\mathcal{R}$ obtained, by aggregating best values for a given instance type and size. We can clearly see the biggest advantage of SBM over SA – its ability to efficiently evolve many replicas simultaneously, while keeping the runtime low. In terms of quality-runtime trade-off, it holds its ground even against the HUBO SA variant, which does not suffer overheads from rank reduction. In particular, SBM is able to, on average, reach solutions of reasonable quality, $\mathcal{R}\approx 0.95$ for Cauchy instances and $\mathcal{R}\approx 0.90$ for Pareto instances, close to an order of magnitude faster than SA, except for large Pareto instances, which seems to be “easy” for the HUBO SA. This is promising, given that it is possible to extend SBM so that it can handle higher-order interactions natively [71]. Finally, insets reveal broad distribution of approximation ratios over random coupling selection for a fixed topology, which further highlights the importance of conducting a proper statistical analysis before drawing any conclusions.
@@
Appendix A4 Simulated Annealing for QUBO and HUBO
@@
  for $pass\in\{1,\,\ldots,\,num\_passes\}$ do
@@
All classical benchmarks in this work were performed on a workstation equipped with dual Intel(R) Xeon(R) Platinum 8462Y+ CPUs, $4$ NVIDIA H100 SXM GPUs, and 11TB of RAM. All results were obtained using a single GPU, unless stated otherwise.